OpenStudy (anonymous):

Use distributive property to write the expression without parentheses: 3(a+b)

OpenStudy (anonymous):

Distributive property, the common variable / constant is "distributed" i.e. multiplied across the bracketed expression. I.e. a(b+c), a is the common, (b+c) is the bracketed expression, i.e. a(b+c) = a(b) + a(c) * a is "distributed" across b and c. 3(a+b) = ?

OpenStudy (anonymous):

So, 3(a+b) = 3(a) +3(b)...is this correct?

OpenStudy (anonymous):

Hello, is this the correct answer??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3a+3b

OpenStudy (anonymous):

TY..is there a way for me to ask graph questions??

OpenStudy (anonymous):

Should it read: 3a +3b? or should it read: 3(a) + 3(b)

OpenStudy (anonymous):

either way is fine you dont really need the parentheses

OpenStudy (anonymous):

ok, I appreciate your help

OpenStudy (anonymous):

mhm :)